I need help with these problems please.
53. Business and finance. The inventor of a new product believes that the cost of producing the product is given by the function
C(x) = 1.75x + 7000
How much does it cost to produce 2000 units of her invention?
54. Business and finance. If the inventor in exercise 53 charges $4 per unit, then her profit for producing and selling x units is given by the function
P(x) = 2.25x - 7000
(a) What is her profit if she sells 2000 units?
(b) What is her profit if she sells 5000 units?
(c) What is the break-even point for sales?
Thanks so much for your help!

Answer by stanbon(48526)

53. Business and finance. The inventor of a new product believes that the cost of producing the product is given by the function
C(x) = 1.75x + 7000
How much does it cost to produce 2000 units of her invention?
C(1000) = 1.75*1000+7000
C(1000) = 1750 + 7000 = $8750
-----------------------------------
54. Business and finance. If the inventor in exercise 53 charges $4 per unit, then her profit for producing and selling x units is given by the function
P(x) = 2.25x - 7000
(a) What is her profit if she sells 2000 units?
P(2000) = 2.25*2000-7000 = 4500 - 7000 = -$2500
-----------------
(b) What is her profit if she sells 5000 units?
P(5000) = 2.25*5000-7000 = 11250-7000 = $4250
--------------------
(c) What is the break-even point for sales?
2.25x-7000=0
2.25x = 7000
x = 3111 or 3112 units will result in no profit or "break even".
==================
Cheers,
Stan H.